OpenStudy (anonymous):

Helen went to the grocery store to buy potatoes. The store charges $1.06 for each pound of potatoes. On the way out of the store, Helen picked up a gallon of ice cream at a cost of $4.32. If Helen spent $12.80 at the store, how many pounds of potatoes did she purchase? Assume tax is included in the given price.

OpenStudy (anonymous):

So the total is 12.80 so you know the potatoes plus the ice cream are going to equal that price. The ice cream is also a constant factor, so you know that won't have any variable. The amount of potatoes, however, does change so that needs a variable. 1.06x+4.32=12.80

OpenStudy (anonymous):

Can you solve it from there? ^^

OpenStudy (anonymous):

Not Really

OpenStudy (anonymous):

1.06x+4.32=12.80 You need to isolate the x, so first subtract 4.32 from both sides and tell me what you get

OpenStudy (anonymous):

A.9 B. 7 C. 8 D.14 Here's the answer choices

OpenStudy (hanamalik):

B 1.06+4.32=5.38 12.80-5.38=7.42 7.42/1.06=7 lbs of potatoes 1.06+4.32+12.80-5.38=12.80 So, Helen left with 7 lbs of potatoes

OpenStudy (anonymous):

Yess! Thanksss

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

How

OpenStudy (anonymous):

1.06+4.32=5.38 says Helen only bought one potato and a thing of ice cream

OpenStudy (hanamalik):

B is the correct answer, just saying.

OpenStudy (anonymous):

If you check your work, 7 potatos would be 7.42 and the icecream is 4.32

OpenStudy (hanamalik):

No variables are needed for this type of equation.

OpenStudy (anonymous):

7.42+4.32 is 11.74

OpenStudy (anonymous):

not 12.80

OpenStudy (anonymous):

The monthly cost of electricity in Smithville is determined by the function C(e) = $0.18e + $27.41, where e is the amount of electricity used in kilowatt-hours (kWh) and C(e) is the total cost. If the Fox family used 853 kWh of electricity last month, how much was their bill?

OpenStudy (anonymous):

1.06x+4.32=12.80 1.06x=8.48 x=8 1.06(8)+4.32 8.48+4.32 12.80 The answer is C

OpenStudy (hanamalik):

$0.18(853)+27.41=$180.95

OpenStudy (anonymous):

The problem says e is the amount of electricty used, and the Fox faily used 853 kWH, so plug 853 where ever theres an E C(e) = $0.18(853) + $27.41

OpenStudy (anonymous):

Oooh

OpenStudy (anonymous):

A company sells each product they make for $90.00. They pay $2,800.00 in monthly costs, such as rent and bills, as well as $75.00 in manufacturing costs per unit. If the company sold 492 units last month, what was their total profit?

OpenStudy (hanamalik):

how many products did they make?

OpenStudy (anonymous):

You need to subtract the price of bills and manufacturing cost per unit from the amount of money they make per unit 90(492)-(2800+75(492)) 44280-(2800+36900) 44280-39700 $4,580.00 profit

OpenStudy (anonymous):

A. $34,100.00 B. $41,405.00 C. $4,580.00 D. $6,056.00

OpenStudy (anonymous):

So It's c

OpenStudy (anonymous):

Yep ^^

OpenStudy (anonymous):

Indicate which property is illustrated in Step 5. Step 1 10 + 9x - 10 - 6x + 5 = 10 + (9x - 10) - 6x + 5 Step 2 = 10 + (-10 + 9x) - 6x + 5 Step 3 = [10 + (-10)] + (9x - 6x) + 5 Step 4 = 0 + (9x - 6x) + 5 Step 5 = (9x - 6x) + 5 Step 6 = (9 - 6)x + 5 Step 7 = 3x + 5

OpenStudy (anonymous):

distributive additive identity associative additive inverse

OpenStudy (anonymous):

In step 5 you got rid of the +0 because adding zero to a number will equal that number so its not really needed. Which one of those four properties deals with adding?

OpenStudy (anonymous):

I Think b

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?7%20 \%20-%20\%20\frac{10}{11}x%20\%20\leq%20\%2017